When doing the f(x)= function, when I get an answer of 3.5625, how do i convert this number into a fraction? I divided this by 100, but my answers aren't correct. I want to be able to convert easily, when I see a problem like this in the future?
Answer by Alan3354(69443)

3.5625, how do i convert this number into a fraction?
--------
= 3 9/16 = 57/16
==============
0.5625 = 5625/10000
Reduce the fraction --> 9/16